OpenStudy (anonymous):

sinxcosxtanx+sinxcosx=0

OpenStudy (anonymous):

\[\sin x \cos x \tan x +\sin x \cos x =0\]

myininaya (myininaya):

factor the sin(x)cos(x) out sin(x)cos(x)[tan(x)+1]=0

myininaya (myininaya):

set all three of those factors equal to 0